# A Complexity-Based Theory of Compositionality

### v1

@@ 2 Compressing a representation
We will argue that a natural way to think about compositional representations is: representations $Z$ that can be significantly compressed as a function of constituent parts. In other words, the shortest program that outputs the representation, with length $K(Z)$, has a very particular form: it first describes $Z$ using short parts-based constituents, and then maps these parts to the high-dimensional representation. This program form is shown in Figure 1 and described in detail below. Crucially, the components of this program will be used in Section 3 to construct our formal definition of compositionality, in which representations that are more compressible as a function of constituent parts are more compositional. Before combining them into a definition of compositionality, we now describe the components of this program in the following steps.
Figure 1: Hypothesized form of the shortest program that outputs a compositional representation $Z$. a. Pseudocode showing the skeleton of the program. The program describes the representation using sentences $W$ (sequences of discrete tokens) that are compressed using a prior $p_{w}(w)$, and then maps these sentences to high-dimensional vectors in representation space using a function $f(w)$ that outputs the sufficient statistics of a Normal distribution. Reconstruction errors are corrected using bit sequences who’s length depends on the magnitudes of the errors. decode_algo() is a short function that decodes an object compressed using arithmetic coding (Witten et al.,, 1987). b. A visual depiction of the program. c. The total Kolmogorov complexity of the representation is estimated by the length of the shortest program that has this form.
@@
Usually, $f(w_{n})$ will not perfectly reconstruct any of the $z_{n}$’s, since $w_{n}$ is discrete and $z_{n}$ is continuous. Since Kolmogorov complexity is about lossless compression, these errors need to be corrected. The number of bits needed for this correction depends on the magnitudes of the errors in the following way. Instead of outputting a vector in $\mathbb{R}^{D}$ representing $z_{n}$ directly, let $f$ output the sufficient statistics of some distribution in $\mathbb{R}^{D}$. At this point, we can evaluate the probability of the true $z_{n}$’s under this distribution, $p(z_{n};f(w_{n}))$, and the number of bits needed to specify $z_{n}$ is equal to $-\log p(z_{n};f(w_{n}))$. For simplicity, we take $p$ to be a Normal distribution who’s mean and standard deviation are given by $f(w_{n})$, in which case the number of correction bits needed depends on the error of $f(w_{n})$: the further the predicted mean is from the true $z_{n}$ and the higher the uncertainty of the prediction, the lower the probability $p(z_{n};f(w_{n}))$ and the higher the correction bit-length $-\log p(z_{n};f(w_{n}))$.
@@ 3 Representational compositionality: a formal definition of compositionality
Crucially, $p_{w}$, $W$, and $f$ are not free parameters: they are intrinsic to the representation in that they best compress $Z$ (see the minimization in Equation 1). Like Kolmogorov complexity, then, $C(Z)$ is intractable to compute because it requires an exponentially-large search over all possible tuples $(p_{w},W,f)$. However, like Kolmogorov complexity, $C(Z)$ can still be tractably estimated using efficient compression and optimization methods. While the primary contribution of this work is theoretical and aimed at justifying Definition 2, we outline a strategy for finding $(p_{w},W,f)$ and estimating $C(Z)$ in Appendix B. Importantly, however, we foreshadow a complementary definition for the compositionality of a language together with a representation, as opposed to a representation alone, which we will introduce in Section 3.1. This quantity is readily estimable in a wide rage of cases, and could prove useful in a number of practical applications such as tokenization.
We now unpack Definition 2 to see how it accounts for the problems of the colloquial Definition 1 and explains computational properties typically associated with compositionality.
Expressivity and compression
Effectively, representational compositionality says that the compositionality of a representation is a compression ratio that depends on two things: (1) the complexity of the representation, which appears in the numerator, and (2) the complexity of the semantics which construct the representation from its constituent parts, which appears in the denominator. When a representation is highly expressive (high $K(Z)$) but can nevertheless be compressed as a simple function of constituent parts (low $K(Z|W)$), representational compositionality says that the representation is highly compositional. Representational compositionality therefore formalizes a hypothesis in cognitive science that compositionality emerges from competing pressures for expressivity and compression (e.g., Kirby,, 1999; Kirby et al.,, 2004, 2008, and references therein).
Constituent “parts” are intrinsic to $Z$
Note that unlike the colloquial Definition 1, representational compositionality makes it clear where the “constituent parts” (tokens in $W$), “complex expressions” ($W$), and “structure” ($f$) associated with a representation come from: they are intrinsic properties of the representation. Compositional representations are those that are compressible in principle as simple functions of constituent parts, regardless of whether or not we know what that optimal compression scheme is.
Systematicity and generalization
Representational compositionality formalizes the intuition that the constituent parts of a compositional representation determine the meaning of the whole in a systematic way (Szabó,, 2022, 2012). For instance, if $f$ arbitrarily maps sentences $w$ to their representations $z$ in a way that does not take the structure or words of the sentence into account (as in the case of idioms), then its complexity $K(f)$ is necessarily high and compositionality is low (we demonstrate this through experiments in Section 4.1). In addition, if $f$ is inaccurate in how it maps sentences to their representations, the error $K(Z|W,f)$ is high and the compositionality low. A representation that is highly compositional according to our definition thus benefits from the generalization ability of simple functions (low $K(f)$) that fit their data well (low $K(Z|W,f)$). Crucially, this ability of $f$ to generalize to novel sentences and representation samples explains the fundamental relationship between compositionality and notions of systematicity from cognitive science (Szabó,, 2022).
Structure-preserving semantics
Representational compositionality explains the widely-held intuition that semantics functions $f$ which are compositional are structure-preserving in how they map $w\rightarrow z$ (Montague et al.,, 1970). As explained in Ren et al., (2023), structure-preserving maps have lower Kolmogorov complexity, and thus higher compositionality according to our definition. In a structure-preserving map, each word in the sentence $w$ independently affects a different subspace of the representation $z$ so that pairwise-distances are similar in sentence-space and representation-space.
Modularity & compositionality
Representational compositionality explains the precise relationship between compositionality and structural modularity, which has been taken for granted in past work but is difficult to formally articulate (Lepori et al.,, 2023; Goyal and Bengio,, 2022; Mittal et al.,, 2022). A modular semantics function $f$ is simple because it decomposes knowledge into smaller reusable components that are algorithmically independent, and thus contributes to high compositionality under our definition. This also explains why natural language is highly compositional. Linguists typically model language using context-free grammars (Chomsky,, 1956), in which a sentence hierarchically decomposes into a parse tree with a “production rule” applied at each node. The recursive application of these production rules, akin to a small number of modules in $f$, is then thought to determine the meaning of the sentence as a whole.
3.1 Compositionality of language systems
In Definition 2 of representational compositionality, $W$ is not a free parameter, but rather a collection of sentences intrinsic to $Z$ that minimize its description length. However, we can also consider the special case of languages in which the sentences are fixed to some $W^{L}$ that is external to the representation. Such cases are common in the real world; in a natural language such as English, $W^{L}$ are the sentences that a person may utter while $Z$ are the underlying neural activity patterns (thoughts) that those sentences elicit. We could then ask to what degree this language system composed of thoughts $Z$ and given sentences $W^{L}$ is compositional. Expressing each thought with an arbitrary sentence would result in a non-compositional language system, whereas using sentences that accurately describe thoughts through simple semantics would be highly compositional. We define the compositionality of a language system as:
@@ 3.1 Compositionality of language systems
This definition opens the door to comparisons between the compositionalities of different real-world language systems, such as French and Japanese, which we attempt in Section 4.3.
@@ 4 Empirical results
While no other formal definition of representational compositionality has been proposed, a heuristic commonly used to measure language system compositionality is topological similarity. For some language system $(W^{L},Z)$, topological similarity computes a pairwise distance matrix between rows in $W^{L}$ and another distance matrix for $Z$, then correlates the two flattened matrices. Intuitively, if the pairwise distances of $Z$ are preserved in $W^{L}$, topological similarity scores the language system as compositional because the two spaces share structure. Throughout our experiments, we compare our definitions to topological similarity. As an aside, we note that our definition explains why topological similarity is a reasonable heuristic: simple semantics functions $f$ (e.g., identity, linear) tend to preserve the structure of their inputs.
@@ 4.1 Synthetic representations
Our first set of experiments consider representations $Z$ that are generated synthetically using known rules through: $z\sim p(z;f(w)),w\sim p_{w}(w)$. Since we know the underlying programs that generated the representations in this case, we know the ground truth complexity terms $K(p_{w})$, $K(W|p_{w})$, $K(f)$, and $K(Z|W,f)$ needed to compute $C(Z)$ exactly. These experiments are therefore less about empirically estimating compositionality in practice and more about validating whether the definition matches with intuitions. The high-level methodology used to generate representations is described below, with additional details (such as derivations of complexity terms) provided in Appendix F.
@@
Vocabulary size: Vocabulary size has a more complex relationship to compositionality. If the vocabulary is too small relative to sentence length, then expressivity and compositionality are limited (e.g., with only one word in the vocabulary, nothing can be expressed). On the other hand, if the vocabulary is too large relative to sentence length, then compositionality is low because expressivity doesn’t come from combining constituent parts (e.g., with one-word sentences and a large vocabulary, there is no notion of parts). For a given sentence length, then, compositionality should peak at some intermediate vocabulary size. This is precisely what we observe empirically with representational compositionality: a sharp increase in compositionality early on followed by a monotonic decrease as vocabulary size increases further. While topological similarity also decreases as a function of increased vocabulary size, it does not show the early increase, and is in fact largest for a vocabulary size of $1$.
@@
To model such systems using representational compositionality, we generated representations using simple synthetic grammars (Figure 2c). First, we assigned each word in the vocabulary an embedding and a part of speech tag, and we defined a grammar with a set of production rules. We then generated a dataset of sentences and parsed them using the grammar. Finally, the semantics were defined by embedding each word in the sentence and then applying a rule-specific function at every node in the parse tree until the root was reached, who’s value we defined to be the representation. The rule-specific functions concatenated children embeddings and applied a linear projection with rule-specific weights.
@@ 4.2 Emergent languages from multi-agent training
Figure 3: Compositionality of language systems that emerge in multi-agent settings with and without iterated learning. a. We consider language systems in which $W^{L}$ and $Z$ emerge from agents trained to communicate in an object reference game. Iterated learning is an inductive bias for compositionality in which periodically resetting model parameters creates simpler languages across generations. We trained models with and without iterated learning to compare their compositionalities using our definition. b. We used prequential coding to measure $K(Z|W^{L})$ for the emergent languages, where the area under the curve is the “prequential code length” estimating compression size. $W^{L}$ for models trained using iterated learning achieved a much lower prequential code length than those trained normally without iterated learning, meaning the semantics $f$ mapping $W^{L}$ to $Z$ were simpler. c. Our language system compositionality metric $C^{L}(Z)$ (blue) agrees with topological similarity (red) on the ordering of models trained with and without iterated learning, but the numerical values provided by $C^{L}(Z)$ provide more theoretical insight (see main text). All error bars show standard deviations across $5$ random seeds.
@@
Next, we further validate our compositionality metric by applying it to real-world representations. To avoid having to solve the difficult optimization problem involved in measuring $C(Z)$ (which requires a minimization of $K(Z)$ w.r.t. $p_{w}$, $W$, $f$) we instead consider language systems in which $W=W^{L}$ is fixed and measure $C^{L}(Z)$ (see Section 3.1).
@@
One interesting case of real language systems is those that emerge in multi-agent settings where agents must learn to communicate. We consider the setting of Li and Bowling, (2019); Ren et al., (2020) in which a speaker and a listener learn to communicate in a simple object reference game, where objects have symbolic attributes analogous to color, size, shape, etc. Agents trained using reinforcement learning typically communicate successfully, but often learn non-compositional language systems that arbitrarily map sentences to objects. However, Li and Bowling, (2019); Ren et al., (2020) have shown that compositionality can emerge through a multi-generation process called iterated learning (Kirby et al.,, 2015), where the agents’ parameters are periodically reset and retrained on sentence/object pairs from the previous generation (see Figure 3a). Kirby et al., (2015) hypothesize that this occurs because iterated learning places an inductive bias for simple language systems that are more easily learnable across subsequent generations.
@@
We trained agents both with and without iterated learning and measured $C^{L}(Z)$ for the resulting language systems. Training details are provided in Appendix G. After $N$ generations, we obtain a dataset consisting of all possible objects $Z$ and the sentences output by the speaker $W^{L}$ when given those objects as input. To measure $C^{L}(Z)$, we need both $K(Z)$ and $K(Z|W^{L})$. Since $Z$ consists of a set of symbolic objects sampled uniformly, $K(Z)$ is simply equal to $|\mathcal{O}|\log_{2}(|\mathcal{O}|)$, where $\mathcal{O}$ is the set of all possible objects. To measure $K(Z|W^{L})$, we used a compression method called prequential coding (Blier and Ollivier,, 2018) that provides good estimates in practice (see Appendix E). Intuitively, prequential coding compresses $Z$ given $W$ by incrementally encoding individual datapoints $z_{<i}$ and fitting a model $\theta_{i-1}$ to predict them using $w_{<i}$ as input. The more datapoints are encoded, the better the model becomes by having seen more training data, and the more accurately it can predict the next datapoint $z_{i}$. Since prediction error is equivalent to complexity, $K(z_{i}|w_{i},\theta_{i-1})$ will decrease as a function of $i$, which means that every subsequent datapoint takes fewer bits to encode. The total complexity $K(Z|W)$ is estimated by summing all of these terms.

### v3

@@ 2 Compressing a representation
We will argue that a natural way to think about compositional representations is: representations $Z$ that can be significantly compressed as a function of constituent parts. In other words, the shortest program that outputs the representation, with length $K(Z)$, has a very particular form: it first describes $Z$ using short parts-based constituents, and then maps these parts to the high-dimensional representation. This program form is shown in Figure 1 and described in detail below. We also give a summary of all program components in Table 1. Crucially, the components of this program will be used in Section 3 to construct our formal definition of compositionality, in which representations that are more compressible as a function of constituent parts are more compositional. Before combining them into a definition of compositionality, we now describe the components of this program in the following steps.
Figure 1: Assumed form of the shortest program that outputs a compositional representation $Z$. a. Pseudocode showing the skeleton of the program. The program describes the representation using sentences $W$ (sequences of discrete tokens) that are compressed using a prior $p_{w}(w)$, and then maps these sentences to high-dimensional vectors in representation space using a function $f(w)$ that outputs the sufficient statistics of a Normal distribution. Reconstruction errors are corrected using bit sequences whose length depends on the magnitudes of the errors. decode() is a short function that decodes an object compressed using arithmetic coding (Witten et al.,, 1987). b. Illustration of the program compressing a representation from a pretrained model layer, brain region, etc. c. The total Kolmogorov complexity of the representation is estimated by the length of the shortest program that has this form.
Table 1: Components of assumed shortest program that outputs a compositional representation $Z$
Name Symbol Example (for representations of scene images)
Representation $Z\in\mathbb{R}^{N\times D}$ Layer activations of a CNN in response to $N$ scene images
Sentences $W\in\mathcal{V}^{N\times M}$ Symbol sequence expressing a scene graph for each $z\in Z$
Language $p_{w}$ Distribution over sentences expressing scene graphs
Semantics $f$ Embed & concatenate each object/relation in the scene graph
Recon. error $\mathcal{N}(z;f(w))$ Correct remaining error unaccounted for by the semantics
@@
Usually, $f(w_{n})$ will not perfectly reconstruct any of the $z_{n}$’s, since $w_{n}$ is discrete and $z_{n}$ is continuous. Since Kolmogorov complexity is about lossless compression, these errors need to be corrected. The number of bits needed for this correction depends on the magnitudes of the errors in the following way. Instead of outputting a vector in $\mathbb{R}^{D}$ representing $z_{n}$ directly, let $f$ output the sufficient statistics of some distribution in $\mathbb{R}^{D}$. At this point, we can evaluate the probability of the true $z_{n}$’s under this distribution, $\mathcal{N}(z_{n};f(w_{n}))$, and the number of bits needed to specify $z_{n}$ is equal to $-\log\mathcal{N}(z_{n};f(w_{n}))$. For simplicity, we take $p$ to be a Normal distribution whose mean and standard deviation are given by $f(w_{n})$, in which case the number of correction bits needed depends on the error of $f(w_{n})$: the further the predicted mean is from the true $z_{n}$ and the higher the uncertainty of the prediction, the lower the probability $\mathcal{N}(z_{n};f(w_{n}))$ and the higher the correction bit-length $-\log\mathcal{N}(z_{n};f(w_{n}))$.
@@ 3 Representational compositionality: a formal definition of compositionality
Crucially, $p_{w}$, $W$, and $f$ are not free parameters: they are intrinsic to the representation in that they best compress $Z$ (see the minimization in Equation 1). Like Kolmogorov complexity, then, $C(Z)$ is intractable to compute because it requires an exponentially-large search over all possible tuples $(p_{w},W,f)$. However, like Kolmogorov complexity, $C(Z)$ can still be tractably estimated using efficient compression and optimization methods. In Appendix B, we outline a strategy for estimating $(p_{w},W,f)$ and thus, $C(Z)$ for general cases, though we leave precise implementation for future work. Importantly, Definition 2 can also be adapted for language systems where a symbolic mapping $W$ (a language) is given and fixed:
@@
Intuitively, $C(Z)$ measures how expressive a representation is relative to how well it can be described as a simple function of parts. The parts $W$ can be thought of as the intrinsic language of the representation, in the sense that this language enables optimal compression. In $C^{L}(Z)$, however, the language $W^{L}$ is externally defined; for instance in a natural language, $W^{L}$ are the sentences that a person might utter while $Z$ are the neural activity patterns (thoughts) that those sentences elicit. Both definitions have complementary value. While $C(Z)$ is more general, $C^{L}(Z)$ can be used to estimate the compositionalities of real-world languages (which we attempt in Section 4.3) or other mappings between symbolic sequences and representations (e.g. tokenization schemes), making it a useful tool for AI development. In addition, $C^{L}(Z)$ can be used to investigate a learned representation’s compositionality with respect to a dataset’s known underlying generative factors, which relates it to much prior work on compositionality (Ren and Sutherland,, 2024; Wiedemer et al.,, 2023; Ren et al.,, 2023). Finally, $C^{L}(Z)$ is easier to estimate since $W^{L}$ is given.
In Section 4, we will first illustrate and test properties of $C(Z)$ using synthetic data where where computations are tractable. We follow with real-world evaluation of $C^{L}(Z)$ on various language systems. Before doing so, we begin by unpacking our definitions to see how they account for the problems of the colloquial Definition 1 and explain computational properties typically associated with compositionality.
Expressivity and compression
Effectively, representational compositionality says that the compositionality of a representation is a compression ratio that depends on two things: (1) the complexity of the representation, which appears in the numerator, and (2) the complexity of the semantics which construct the representation from its constituent parts, which appears in the denominator. When a representation is highly expressive (high $K(Z)$) but can nevertheless be compressed as a simple function of constituent parts (low $K(Z|W)$), representational compositionality says that the representation is highly compositional. Representational compositionality therefore formalizes a hypothesis in cognitive science that compositionality emerges from competing pressures for expressivity and compression (e.g., Kirby,, 1999; Kirby et al.,, 2004, 2008, and references therein), which has also recently been explored in AI from a theoretical perspective (Ren and Sutherland,, 2024)..
Constituent “parts” are intrinsic to $Z$
Note that unlike the colloquial Definition 1, representational compositionality makes it clear where the “constituent parts” (tokens in $W$), “complex expressions” ($W$), and “structure” ($f$) associated with a representation come from: they are intrinsic properties of the representation defined through the lens of optimal compression. Compositional representations are those that are compressible in principle as simple functions of constituent parts, regardless of whether or not we know what that optimal compression scheme is. This is a significant difference between Definition 2 and other related ideas in the literature which quantify compositionality in terms of reconstruction from externally-defined parts (e.g., Andreas,, 2019; Trager et al.,, 2023; Lewis et al.,, 2022). In addition, unlike prior (Andreas,, 2019; Ram et al.,, 2024), our definition makes no strong assumptions about the form of the reconstruction (e.g., that it is linear, a hierarchical grammar, etc.) as it abstracts over arbitrary functions through the lens of their complexity $K(f)$. Definition 2 therefore generalizes diverse notions of compositionality framed in terms of representation-reconstruction.
Systematicity and generalization
Representational compositionality formalizes the intuition that the constituent parts of a compositional representation determine the meaning of the whole in a systematic way (Szabó,, 2022, 2012), where “systematicity” is a term from cognitive science that roughly means “structured” or “non-arbitrary”. If $f$ arbitrarily maps sentences $w$ to their representations $z$ in a way that does not take the structure or words of the sentence into account (as in the case of idioms), then its complexity $K(f)$ is necessarily high and compositionality is low (we demonstrate this through experiments in Section 4.1). In addition, if $f$ is inaccurate in how it maps sentences to their representations, the error $K(Z|W,f)$ is high and the compositionality low. A representation that is highly compositional according to our definition thus benefits from the generalization ability of simple functions (low $K(f)$) that fit their data well (low $K(Z|W,f)$). Crucially, this ability of $f$ to generalize to novel sentences and representation samples explains the fundamental relationship between compositionality and notions of systematicity from cognitive science (Szabó,, 2022).
Structure-preserving semantics
Representational compositionality explains the widely-held intuition that semantics functions $f$ which are compositional are structure-preserving in how they map $w\rightarrow z$ (Montague et al.,, 1970). In a structure-preserving map, each word in the sentence $w$ independently affects a different subspace of the representation $z$ so that pairwise-distances are similar in sentence-space and representation-space. As explained in Ren et al., (2023), structure-preserving maps have lower Kolmogorov complexity, and thus higher compositionality according to our definition.
Modularity and compositionality
Representational compositionality explains the precise relationship between compositionality and structural modularity, which has been taken for granted in past work but is difficult to formally articulate (Lepori et al.,, 2023; Goyal and Bengio,, 2022; Mittal et al.,, 2022). A modular semantics function $f$ is simple because it decomposes knowledge into smaller reusable components that are algorithmically independent and only need to be defined once, thus contributing to high compositionality under our definition. This also explains why natural language is highly compositional. Linguists typically model language using context-free grammars (Chomsky,, 1956), in which a sentence hierarchically decomposes into a parse tree with a “production rule” applied at each node. The recursive application of these production rules, akin to a small number of modules in $f$, is then thought to determine the meaning of the sentence as a whole.
@@ 4 Empirical results
While no other formal definition of representational compositionality has been proposed, a heuristic commonly used to measure language system compositionality is topological similarity. For some $(W,Z)$, topological similarity computes a distance between all pairs of sentences $\Delta_{\mathcal{W}}^{ij}=d_{\mathcal{W}}(w_{i},w_{j})$ using a distance metric $d_{\mathcal{W}}(\cdot)$ in $\mathcal{W}$, and a distance between all pairs of representation elements $\Delta_{\mathcal{Z}}^{ij}=d_{\mathcal{Z}}(z_{i},z_{j})$ using a distance metric $d_{\mathcal{Z}}(\cdot)$ in $\mathcal{Z}$. It then computes the Pearson correlation $\rho$ between the two pairwise distance matrices, quantifying the degree to which the two spaces share linear structure. Throughout our experiments, we compare our definitions to topological similarity. Throughout our experiments, we compare our definitions to topological similarity. As an aside, we note that our definition explains why topological similarity is a reasonable heuristic: simple semantics functions $f$ (e.g., identity, linear) tend to preserve the structure of their inputs.
@@ 4.1 Synthetic representations
Our first set of experiments consider representations $Z$ that are generated synthetically using known rules through: $z\sim\mathcal{N}(z;f(w)),w\sim p_{w}(w)$. Since we know the underlying programs that generated the representations in this case, we know the ground truth complexity terms $K(p_{w})$, $K(W|p_{w})$, $K(f)$, and $K(Z|W,f)$ needed to compute $C(Z)$ exactly. These experiments are therefore less about empirically estimating compositionality in practice and more about validating whether the definition matches with intuitions. The high-level methodology used to generate representations is described below, with additional details (such as derivations of complexity terms) provided in Appendix H.
@@
Vocabulary size: Vocabulary size has a more complex relationship to compositionality. If the vocabulary is too small relative to sentence length, then expressivity and compositionality are limited (e.g., with only one word in the vocabulary, nothing can be expressed). On the other hand, if the vocabulary is too large relative to sentence length, then compositionality is low because expressivity doesn’t come from combining constituent parts (e.g., with one-word sentences and a large vocabulary, there is no notion of parts). For a given sentence length, then, compositionality should peak at some intermediate vocabulary size. This is precisely what we observe empirically with representational compositionality: a sharp early increase in compositionality followed by a monotonic decrease as vocabulary size increases further. While topological similarity also decreases as a function of increased vocabulary size, it does not show the early increase, and is in fact largest for a vocabulary size of $1$.
@@
To model such systems using representational compositionality, we generated representations using simple synthetic grammars (Figure 2c). First, we assigned each word in the vocabulary an embedding and a part of speech tag, and we defined a grammar with a set of production rules. We then generated a dataset of sentences and parsed them using the grammar. Finally, the semantics were defined by embedding each word in the sentence and then applying a rule-specific function at every node in the parse tree until the root was reached, whose value we defined to be the representation. The rule-specific functions concatenated children embeddings and applied a linear projection with rule-specific weights.
@@ 4.2 Emergent languages from multi-agent training
Figure 3: Compositionality of language systems that emerge in multi-agent settings with and without iterated learning. a. We consider language systems in which $W^{L}$ and $Z$ emerge from agents trained to communicate in an object reference game. Iterated learning amplifies a model’s inductive bias for learning compositional languages more quickly, where periodically resetting model parameters creates simpler languages across generations. We trained models with and without iterated learning to compare their compositionalities using our definition. b. We used prequential coding to measure $K(Z|W^{L})$ for the emergent languages, where the area under the curve is the “prequential code length” estimating compression size. $W^{L}$ for models trained using iterated learning achieved a much lower prequential code length than those trained normally without iterated learning, meaning the semantics $f$ mapping $W^{L}$ to $Z$ were simpler. c. Our language system compositionality metric $C^{L}(Z)$ (blue) agrees with topological similarity (red) on the ordering of models trained with and without iterated learning, but the numerical values provided by $C^{L}(Z)$ provide more theoretical insight (see main text). All error bars show standard deviations across $5$ random seeds.
@@
Next, we further validate our compositionality metric by applying it to real-world representations. To avoid having to solve the difficult optimization problem involved in measuring $C(Z)$ (which requires a minimization of $K(Z)$ w.r.t. $p_{w}$, $W$, $f$) we instead consider language systems in which $W=W^{L}$ is fixed and measure $C^{L}(Z)$ through Definition 3.
@@
One interesting case of real language systems is those that emerge in multi-agent settings where agents must learn to communicate. We consider the setting of Li and Bowling, (2019); Ren et al., (2020) in which a speaker and a listener learn to communicate in a simple object reference game, where objects have symbolic attributes analogous to color, size, shape, etc. Agents trained using reinforcement learning typically communicate successfully, but often learn non-compositional language systems that arbitrarily map sentences to objects. However, Li and Bowling, (2019); Ren et al., (2020) have shown that compositionality can emerge through a multi-generation process called iterated learning (Kirby et al.,, 2015), where the agents’ parameters are periodically reset and pretrained on sentence/object pairs from the previous generation (see Figure 3a). Kirby et al., (2015) hypothesize that this occurs because iterated learning places an inductive bias for simple language systems that are more easily learnable across subsequent generations.
@@
We trained agents both with and without iterated learning and measured $C^{L}(Z)$ for the resulting language systems. Training details are provided in Appendix I. After $N$ generations, we obtain a dataset consisting of all possible objects $Z$ and the sentences output by the speaker $W^{L}$ when given those objects as input. To measure $C^{L}(Z)$, we need both $K(Z)$ and $K(Z|W^{L})$. Since $Z$ consists of a set of symbolic objects sampled uniformly, $K(Z)$ is simply equal to $|\mathcal{O}|\log_{2}(|\mathcal{O}|)$, where $\mathcal{O}$ is the set of all possible objects. To measure $K(Z|W^{L})$, we used a compression method called prequential coding (Blier and Ollivier,, 2018) that provides good estimates in practice (see Appendix G). Intuitively, prequential coding compresses $Z$ given $W$ by incrementally encoding individual datapoints $z_{<i}$ and fitting a model $\theta_{i-1}$ to predict them using $w_{<i}$ as input. The more datapoints are encoded, the better the model becomes by having seen more training data, and the more accurately it can predict the next datapoint $z_{i}$. Since prediction error is equivalent to complexity, $K(z_{i}|w_{i},\theta_{i-1})$ will decrease as a function of $i$, which means that every subsequent datapoint takes fewer bits to encode. The total complexity $K(Z|W)$ is estimated by summing all of these terms.